OpenStudy (anonymous):

Use the trigonometric functions to answer: Find the length of ZX and XY for the triangle XYZ given that: X = 90° Z = 60° ZY = 20mm

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay so you see, that ZX is the base, ZY is the hypotenuse and XY is the perpendicular. Now you must think what trigonometric functions to use to obtain ZX and XY. For ZX, \[\cos \theta= \frac{ base }{ hyp}\] \[\cos 60= \frac{ ZX }{ 20 }\]

OpenStudy (anonymous):

For XY, \[\sin \theta= \frac{ perpendicular }{ hyp }\] \[\sin 60= \frac{ XY }{ 20 }\]

OpenStudy (anonymous):

I think you can continue from here and do the calculations yourself.

OpenStudy (anonymous):

Thank you I wasn't sure about what formulas to use for this.

OpenStudy (anonymous):

I got 17.32 (2dp) for XY. Then for ZX I got 10.

OpenStudy (anonymous):

Yes, your answers are correct. (:

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw.